### v1

# Tunable Molecular Interactions Near an Atomic Feshbach Resonance: Stability and Collapse of a Molecular Bose-Einstein Condensate

Abstract
Understanding and controlling interactions of ultracold molecules has been a central goal in quantum chemistry research. Recent experiments on atoms near a Feshbach resonance offer the key to prepare and investigate molecules in the quantum many-body regime. Just as Feshbach resonances allow tuning of the scattering length of bosonic atoms, we show that they also modify the scattering length of Feshbach molecules which are constituted from these atoms. Based on calculations of the compressibility, we determine the stability phase diagrams of molecular condensates and show that their instability can be associated with a sign change of the inter-molecular interactions. We derive universal expressions for the molecular scattering lengths, presented in terms of experimentally measurable quantities. These will enable control of interactions between Feshbach molecules as well as further studies of few- and many-body reactions involving Feshbach molecules in the quantum regime.
Introduction — Molecular condensates formed from bosonic atoms present exciting opportunities for revealing new states of matter and novel quantum phase transitions Radzihovsky et al. (2004); Romans et al. (2004); Radzihovsky et al. (2008); Radzihovsky and Choi (2009). A key technique for producing these diatomic molecules is through Feshbach resonances Chin et al. (2010), where magnetoassociation converts pairs of ultracold atoms into bound diatomic molecules. This process, involving a sweep of a magnetic field across a resonance, has enabled the preparation of stable molecular condensates in equilibrium Zhang et al. (2021, 2023). Such condensates in more direct proximity to resonance are inaccessible due to severe three-body recombination Fedichev et al. (1996); Bedaque et al. (2000); Weber et al. (2003); Fletcher et al. (2013) and associated particle loss and heating. This situation is in striking contrast to the two-component Fermi gas, where, due to the Pauli exclusion, stable molecular condensates  Petrov et al. (2004, 2005), and as a result the associated BCS-BEC crossover, have been extensively studied Zwerger (2011); Regal et al. (2004); Zwierlein et al. (2004); Chen et al. (2005).
The success reported in Refs. Zhang et al. (2021, 2023) then raises the natural question of how to understand and characterize the stability or, more typically, lack thereof, in these near-resonant molecular condensates  Jeon et al. (2002); Koetsier et al. (2009). In this paper we address this important issue, focusing on the instability of molecular condensates as represented by the behavior of the molecule-molecule scattering length. This is inspired by the behavior of atomic condensates, whose two-body $s$-wave scattering length $a_{s}$ can generally be expressed as
$a_{s}$ $=a_{\mathrm{bg}}\big{(}1-\frac{\Delta B}{B-B_{0}}\big{)},$ (1)
where $B_{0}$ is the resonance value of the magnetic field $B$, $\Delta B$ is the resonance width and $a_{\mathrm{bg}}$ is the atomic $s$-wave background scattering length. The leading-order contribution to $a_{s}$ due to the Feshbach coupling can be represented by Fig. 1(a)  [^1] Higher-order contributions involve multiple repetitions of the process in Fig. 1(a). While these contributions do not alter the resonant structure of $a_{s}$ in Eq. (1),they do lead to a renormalization of $B_{0}$. . When $a_{s}$ is negative, the atomic condensates undergo a collapse Kagan et al. (1998), which occurs when the magnetic field $B$ is above the resonance $B_{0}$ by no more than the resonance width $\Delta B$, or $B_{0}<B<B_{0}+\Delta B$ (for a positive atom background scattering length $a_{\mathrm{bg}}>0$). For concreteness, here we have assumed $\Delta B>0$ so that the energy of the bare molecules falls below that of the atom continuum when the magnetic field is below the resonance $B<B_{0}$.   [^2] The formulae derived here can be easily generalized to the case of negative $\Delta B$.. What has not been widely appreciated is that the Feshbach coupling also resonantly modifies the scattering properties of molecules and can lead to an instability of a molecular Bose-Einstein condensate.
Figure 1: Scattering between atoms (a) and molecules (b) near an atomic Feshbach resonance. (a) Leading-order scattering process that contributes to the resonant term appearing in $a_{s}$ in Eq. (1). This involves two incoming atoms (single blue circles) that temporarily combine into one molecule (two attached circles) through the Feshbach coupling and then dissociate into two outgoing free atoms. The orange region represents virtual processes. (b) The corresponding leading-order contribution to the molecular scattering length $a_{\mathrm{mm}}$ near the resonance. Here, two molecules approach each other, temporarily break up into four free atoms which propagate, interact, and then recombine into another two molecules. As in (a), this process universally depends on the Feshbach coupling, but is of higher order in the coupling strength.
In this paper, on the basis of a two-channel model, we demonstrate that, similar to the atomic condensate case, Feshbach resonances also alter the inter-molecular scattering length $a_{\mathrm{mm}}$. This can lead to the instability of molecular condensates when the molecular scattering length is tuned to a negative value. Importantly, this molecular scattering length can be further modified due to finite density effects which we characterize through an additional effective scattering length $a_{\mathrm{mm}}^{\mathrm{eff}}$. This effective scattering length will depend on the density significantly for a narrow resonance due to many-body effects. In the case of a very wide Feshbach resonance this dependence is weak and can be safely ignored, so that $a_{\mathrm{mm}}^{\mathrm{eff}}$ becomes equivalent to the two-body molecular scattering length $a_{\mathrm{mm}}$.
Arriving at the scattering lengths $a_{\mathrm{mm}}^{\mathrm{eff}}$ and $a_{\mathrm{mm}}$ presents a significant challenge for theory. Hints of this difficulty can already be seen from Fig. 1(b), which indicates just one of the leading order processes which is already of higher order in the coupling strength than the simpler atomic analogue in Fig. 1(a). Additional challenges come from the need to include contributions from background interactions among atoms and molecules.
Summary of main results — Our calculations are based on a variational treatment, from which we address the behavior of the compressibility and deduce the associated scattering lengths. We first summarize some general results.
Figure 2: Phase diagrams and molecular scattering lengths. (a,b) Ground-state stability phase diagrams for (a) wide and (b) narrow resonances. Plotted in (a,b) is a map of the compressibility $\kappa=dn/d\mu$ as a function of atom number density $n$ and the detuning $\bar{\nu}=\Delta\mu_{m}(B-B_{0})$. The compressibility $\kappa$ and the detuning $\bar{\nu}$ are normalized by $\kappa_{\mathrm{bg}}=m_{1}/(4\pi\hbar^{2}a_{\mathrm{bg}})$ and the resonance width $\Delta_{\bar{\nu}}=\Delta\mu_{m}\Delta B$, respectively. The atomic condensate is nonzero only in the ASF phase, while the molecule condensate is nonzero in both the ASF and MSF phases. Stable regions with positive compressibility $\kappa>0$ are shaded in orange, while the unstable region is in blue. The red dashed lines in (a,b) correspond to the quantum critical point at $\bar{\nu}_{c}(n)$, separating the ASF from the MSF phase, and are determined independently of stability. The parameters $\bar{\nu}_{c,-}$ and $\bar{\nu}_{c,+}$ represent the detuning separating the unstable and stable regions. (c,d) Corresponding two-body scattering length $a_{\mathrm{mm}}$, and its many-body analogue $a_{\mathrm{mm}}^{\mathrm{eff}}(n)$ at a typical density $na_{\mathrm{bg}}^{3}=1.68\times 10^{-5}$. For the narrow resonances in (b,d) we choose parameters from the resonance of 133Cs at $B_{0}=19.849$ G used in Refs. Zhang et al. (2021, 2023); Wang et al. (2024), while for the wide resonances in (a,c) we increased the resonance width $\Delta_{\bar{\nu}}$ from its value in (b,d) by $10^{2}$ times, keeping all other parameters the same.
Plotted in Fig. 2(a,b) are the zero-temperature stability phase diagrams for both a wide and narrow resonance, in terms of density $n$ on the vertical axis and detuning $\bar{\nu}$ on the horizontal axis. These plots are obtained by numerical calculations of the compressibility $\kappa$. Both panels indicate a stable molecular superfluid (MSF) phase at large negative detuning, a stable atomic superfluid (ASF) phase at large positive detuning, and an unstable region in between, near the resonance. The instability corresponds to parameters where the compressibility is negative $\kappa<0$. We call attention to the detuning labels $\bar{\nu}_{c,+}$ and $\bar{\nu}_{c,-}$ which delineate the two end points for the unstable region. In Fig. 2(c,d), we plot both our calculated two-body scattering length, $a_{\mathrm{mm}}$, and its many body analogue $a_{\mathrm{mm}}^{\mathrm{eff}}(n)$.
There is a pronounced similarity between panels (a) and (b) of Fig. 2 on the atomic side, as for both narrow and wide resonances, the phase boundary of the unstable region is essentially governed by the sign change of the two-body atomic scattering length $a_{s}$. Here, in both cases the upper boundary detuning $\bar{\nu}_{c,+}$ of the unstable regime is, thus, nearly $n$-independent.
By contrast, the behavior of the compressibility on the molecular side of the phase diagrams in Fig. 2 displays a notable contrast between the narrow and wide resonance cases. We summarize some of our key observations as follows:
  1. 1.
    In the wide resonance case the entire region of instability shows very little density dependence. For this reason we should be able to characterize the molecular-condensate stability entirely in terms of a two body scattering length (that is appropriate to zero density), as is similar to the atomic condensate case. This can be seen from Fig. 2(c) which shows the near-equivalence $a_{\mathrm{mm}}^{\mathrm{eff}}(n)\approx a_{\mathrm{mm}}$ and also the fact that the two-body scattering length $a_{\mathrm{mm}}$ changes sign at the instability onset detuning $\bar{\nu}_{c,-}$ in Fig. 2(a).
  2. 2.
    By contrast, in the narrow resonance case, the lower boundary detuning $\bar{\nu}_{c,-}$ of the unstable region in Fig. 2(b) shows a strong dependence on density. This lower boundary detuning is also the point at which the density-dependent many-body scattering length $a_{\mathrm{mm}}^{\mathrm{eff}}$, rather than the two-body scattering length $a_{\mathrm{mm}}$, changes sign, as seen from Fig. 2(d).
  3. 3.
    Interestingly, unlike in the wide resonance case, the boundary detuning $\bar{\nu}_{c,-}$ in Fig. 2(b) at moderate and high densities nearly coincides with the detuning $\bar{\nu}_{c}(n)$, which is associated with a quantum critical point (QCP) Radzihovsky et al. (2004); Romans et al. (2004) that separates the MSF from the ASF phase. This implies that proximity to the QCP plays a crucial role in the instability of these molecular superfluids.
To quantitatively characterize the observations above, we derive the following three limiting forms for the scattering lengths $a_{\mathrm{mm}}$ and $a_{\mathrm{mm}}^{\mathrm{eff}}$. These formulae are appropriate in the interesting detuning regions where the compressibility changes its sign sup and they are applicable only for the magnetic field below the resonance $B<B_{0}$.
For a wide resonance at large negative detuning, we find that the scattering lengths $a_{\mathrm{mm}}^{\mathrm{eff}}$ and $a_{\mathrm{mm}}$ are given by
$a_{\mathrm{mm}}^{\mathrm{eff}}(n)$ $\approx a_{\mathrm{mm}}=a_{\mathrm{mm}}^{\mathrm{bg}}\big{[}1-\big{(}\frac{\Delta_{B}^{\mathrm{wide}}}{B_{0}-B}\big{)}^{2}\big{]}$ (2a)
where $\text{where }\qquad\Delta_{B}^{\mathrm{wide}}$ $=\frac{\pi}{\sqrt{6}}\sqrt{\frac{\hbar^{2}}{m_{1}(\bar{a}-a_{\mathrm{bg}})^{2}\Delta\mu_{m}}\Delta B}$ (2b)
is an effective molecular resonance width. Here, $a_{\mathrm{mm}}^{\mathrm{bg}}$, which is chosen to be positive  [^3] In this paper, we consider only the case where both $a_{\mathrm{bg}}$ and $a_{\mathrm{mm}}^{\mathrm{bg}}$ are positive., $m_{1}$, and $\Delta\mu_{m}$ are the molecular $s$-wave background scattering length, the atomic mass, and the magnetic moment difference between the open and closed channels, respectively. As in the literature  Lange et al. (2009); Chin et al. (2010) $\bar{a}$ in Eq. (2b) can be taken as $0.96R_{\mathrm{vdW}}$ where $R_{\mathrm{vdW}}$ is the van der Waals length.
For the narrow resonance case at zero density and with detuning somewhat away from unitarity, we can define a two-body molecular scattering length
$a_{\mathrm{mm}}$ $=a_{\mathrm{mm}}^{\mathrm{bg}}\big{[}1-\big{(}\frac{\Delta_{B}^{\text{narrow}}}{B_{0}-B}\big{)}^{\frac{3}{2}}\big{]}$ (3a)
where $\text{where }\quad\Delta_{B}^{\text{narrow}}$ $=\big{(}\frac{a_{\mathrm{bg}}}{a_{\mathrm{mm}}^{\mathrm{bg}}}\big{)}^{\frac{2}{3}}\big{(}\frac{m_{1}a_{\mathrm{bg}}^{2}\Delta\mu_{m}}{\hbar^{2}}\big{)}^{\frac{1}{3}}\;(\Delta B)^{\frac{4}{3}}$ (3b)
is the counterpart molecular resonance width. The scattering process underlying the resonant term in Eq. (3a)  [^4] Note that the whole contribution of the resonance term in Eq. (3a) does not depend on $a_{\mathrm{mm}}^{\mathrm{bg}}$, since the overall prefactor of $a_{\mathrm{mm}}^{\mathrm{bg}}$ is cancelled out by a factor of $1/a_{\mathrm{mm}}^{\mathrm{bg}}$ from $(\Delta_{B}^{\mathrm{narrow}})^{2}$. was illustrated in Fig. 1(b), which involves the exchange of two bosonic atom constituents between molecules. Importantly, the negative sign arises from Bose statistics. It can be shown that for a fermionic Feshbach resonance, $a_{\mathrm{mm}}$ takes a similar form to Eq. (3a), but with the negative sign in front of the resonant term replaced by a positive sign (See Refs. Gurarie and Radzihovsky (2007); Levinsen and Petrov (2011)  [^5] On the other hand, in the fermionic studies of Refs. Gurarie and Radzihovsky (2007); Levinsen and Petrov (2011), the effect of the molecular background interaction $a_{\mathrm{mm}}^{\mathrm{bg}}$ on $a_{\mathrm{mm}}$ was not considered. For additional discussions, see Ref. sup .).
Finally we present an expression for the many-body molecular scattering length $a_{\mathrm{mm}}^{\mathrm{eff}}(n)$ in the case of a narrow resonance at finite density,
$a_{\mathrm{mm}}^{\mathrm{eff}}(n)$ $\approx a_{\mathrm{mm}}^{\mathrm{bg}}\big{[}1-\big{(}\frac{\Delta_{B}^{\text{narrow}}}{B_{0}-B}\big{)}^{\frac{3}{2}}f(B,n)\big{]}$ (4a)
with $\text{with }\;f(B,n)$ $=1+\frac{2\sqrt{2}}{\pi}\ln\big{(}\frac{\bar{\nu}^{2}}{\bar{\nu}^{2}-\bar{\nu}_{c}^{2}(n)}\big{)},$ (4b)
where $\bar{\nu}=\Delta\mu_{m}(B-B_{0})$ is the detuning and $\bar{\nu}_{c}(n)\approx-2\bar{\alpha}\sqrt{n}$ is its corresponding QCP value, with $\bar{\alpha}$ representing the renormalized Feshbach coupling strength defined in Eq. (6) below. Eq. (4) is applicable for the detuning less than the QCP value, $\bar{\nu}<\bar{\nu}_{c}<0$. At zero density, the QCP coincides with the resonance $\bar{\nu}_{c}=0$ and Eq. (4a) reduces to Eq. (3) for the two-body scattering length $a_{\mathrm{mm}}$.
Qualitatively one can understand why many-body physics plays an important role for the molecular condensate stability, as the main fluctuations that destabilize the MSF phase in Fig. 2 are atomic Bogoliubov excitations arising from the presence of the ground-state molecular condensate. These excitations exhibit a significant energy gap and behave like free atoms when the detuning is large and negative, as is relevant in a wide resonance. However, they become soft with a phonon-like, low-energy dispersion as the detuning approaches the QCP. For a sufficiently narrow resonance at a given density, it is this softness of the Bogoliubov excitations, rather than proximity to the two-body resonance point, that drives the instability.
These many-body effects are manifested in the logarithmic term in $f(B,n)$ appearing in Eq. 4. This logarithmic dependence has the important consequence that the boundary detuning $\bar{\nu}_{c,-}(n)$ for the narrow resonance is generally exponentially close to the QCP sup . Importantly, this narrow resonance case presents an interesting opportunity as it provides experimental access to the associated QCP physics from the stable MSF side.
Variational analysis — To arrive at all of these results, we use a variational ground state analysis that incorporates both atomic and molecular condensates, as well as depletion and atomic Cooper-pairing contributions. Some aspects of these stability issues due to the presence of Cooper pairing of bosons have been discussed in prior work Evans and Imry (1969); Nozieres and Saint James (1982); Jeon et al. (2002); Koetsier et al. (2009); Lee and Lee (2010); Basu and Mueller (2008). There the emphasis was primarily restricted to single-channel models.
We adopt a two-channel model Hamiltonian that includes both inter-channel Feshbach coupling and intra-channel background interactions:
$\hat{H}$ $=\sum_{\sigma=1}^{2}\sum_{\mathbf{k}}h_{\sigma\mathbf{k}}a_{\sigma\mathbf{k}}^{\dagger}a_{\sigma\mathbf{k}}-\frac{\alpha}{\sqrt{V}}\sum_{\mathbf{k}_{1},\mathbf{k}_{2}}\big{(}a^{\dagger}_{1\mathbf{k}_{1}}a^{\dagger}_{1\mathbf{k}_{2}}a_{2,\mathbf{k}_{1}+\mathbf{k}_{2}}$
$+h.c.\big{)}+\sum_{\sigma=1}^{2}\sum_{\mathbf{k}_{i}}\frac{g_{\sigma}}{2V}a^{\dagger}_{\sigma\mathbf{k}_{1}}a^{\dagger}_{\sigma\mathbf{k}_{2}}a_{\sigma\mathbf{k}_{3}}a_{\sigma,\mathbf{k}_{1}+\mathbf{k}_{2}-\mathbf{k}_{3}}.$ (5)
Here, $a_{\sigma\mathbf{k}}$ is an annihilation operator for the open-channel atoms (with $\sigma=1$) or closed-channel molecules ($\sigma=2$). The summation $V^{-1}\sum_{\mathbf{k}}$, with $V$ the volume, represents an integral over the momentum $\mathbf{k}$ with a cutoff $\Lambda$, which is needed to regularize ultraviolet divergences in various $\mathbf{k}$ dependent integrals. Finally, we assume three-dimensional isotropy and ignore trap effects.
In Eq. (Tunable Molecular Interactions Near an Atomic Feshbach Resonance: Stability and Collapse of a Molecular Bose-Einstein Condensate), we define kinetic energy contributions $h_{1\mathbf{k}}=(\hbar\mathbf{k})^{2}/2m_{1}-\mu$ and $h_{2\mathbf{k}}=(\hbar\mathbf{k})^{2}/2m_{2}-(2\mu-\nu)$ with $m_{2}=2m_{1}$, $\mu$ the chemical potential, and $\nu$ the bare molecule detuning. The parameter $g_{\sigma}>0$ corresponds to a repulsive intra-channel density-density interaction, and $\alpha$ represents the bare Feshbach coupling between the two channels. The parameters $\{\nu,\alpha,g_{1},g_{2}\}$ depend on the cutoff $\Lambda$ and are related Wang et al. (2024); sup to experimental observables by $\nu=\bar{\nu}+\sqrt{2}\beta\alpha\bar{\alpha}$, $\alpha=\bar{\alpha}\Gamma/\sqrt{2}$, $g_{1}=\bar{g}_{1}\Gamma$, and $g_{2}=\bar{g}_{2}/(1-(2/\pi)\Lambda a_{\mathrm{mm}}^{\mathrm{bg}})$ with $\bar{\alpha}=\sqrt{4\pi\hbar^{2}a_{\mathrm{bg}}\Delta\mu_{m}\Delta B/m_{1}}$, $\bar{g}_{1}=4\pi\hbar^{2}a_{\mathrm{bg}}/m_{1}>0$, $\bar{g}_{2}=4\pi\hbar^{2}a_{\mathrm{mm}}^{\mathrm{bg}}/m_{2}>0$, $\beta=m_{1}\Lambda/(2\pi^{2}\hbar^{2})$, and $\Gamma=1/(1-\beta\bar{g}_{1})$. These relations are chosen to reproduce the atomic scattering length $a_{s}$ in Eq. (1) in the two-atom scattering limit. Finally, the momentum cutoff $\Lambda$ is related to the length scale $\bar{a}$ in Eqs. (2),(3) and (4) by $\Lambda=\pi/(2\bar{a})$.
From the renormalized Feshbach coupling parameter $\bar{\alpha}$ one can define a characteristic length scale $r_{*}$ Gurarie and Radzihovsky (2007); Levinsen and Petrov (2011); Chin et al. (2010); Ho et al. (2012) as
$r_{*}$ $\equiv 4\pi\hbar^{4}/m_{1}^{2}\bar{\alpha}^{2}=\hbar^{2}/a_{\mathrm{bg}}m_{1}\Delta\mu_{m}\Delta B,$ (6)
which allows us to classify the resonance width quantitatively. If the ratio $w_{\mathrm{res}}\equiv a_{\mathrm{bg}}/r_{*}$ is much smaller than unity $w_{\mathrm{res}}\ll 1$, the resonance is viewed as narrow; otherwise, it is classified as wide. For the narrow and wide resonances used in this paper, we choose $w_{\mathrm{res}}\approx 0.006$ Wang et al. (2024); sup and $w_{\mathrm{res}}\sim 1$, respectively   [^6] We note that using an even wider resonance does not change our conclusions. We use the moderately wide resonance to make our numerical calculation easier..
To address the ground-state stability at zero temperature we adopt the following many-body variational wavefunction as an approximation to the true ground state of $\hat{H}$,
$|\Psi_{\mathrm{var}}\rangle$ $=\mathcal{N}^{-1}e^{\sum_{\sigma}\Psi_{\sigma 0}\sqrt{V}a^{\dagger}_{\sigma 0}+\sum_{\mathbf{k}}^{\prime}\sum_{\sigma}\chi_{\sigma\mathbf{k}}\;a^{\dagger}_{\sigma\mathbf{k}}a^{\dagger}_{\sigma-\mathbf{k}}}|0\rangle,$ (7)
where $\{\Psi_{\sigma 0},\chi_{\sigma\mathbf{k}}\}$ are the variational parameters and $\mathcal{N}$ is the normalization factor. In the exponent, the $\mathbf{k}-$sum is over half of $\mathbf{k}-$space, and the prime in the $\mathbf{k}$-summation implies the origin $\mathbf{k}=0$ is excluded. The vacuum $|0\rangle$ satisfies $a_{\sigma\mathbf{k}}|0\rangle=0$ by all annihilation operators $a_{\sigma\mathbf{k}}$. In the spirit of generalized Bogoliubov theory, this variational wavefunction includes only pair-wise correlations between atoms or between molecules in the exponent, which is sufficient for our focus on the detuning regime that is not too close to the resonance Wang et al. (2025).
The trial ground state energy associated with $|\Psi_{\mathrm{var}}\rangle$ is then
$\Omega[\Psi_{10},\Psi_{20},\chi_{1\mathbf{k}},\chi_{2\mathbf{k}}]=\langle\Psi_{\mathrm{var}}|\hat{H}|\Psi_{\mathrm{var}}\rangle,$ (8)
which is a functional of the parameters $\{\Psi_{10},\Psi_{20},\chi_{1\mathbf{k}},\chi_{2\mathbf{k}}\}$. Here, $\Psi_{\sigma 0}=\langle a_{\sigma 0}\rangle/\sqrt{V}$ indicates that $\Psi_{10}$ ($\Psi_{20}$) also represents the amplitude of the atomic (molecular) condensate. Minimizing $\Omega$ with respect to $\{\Psi_{10}^{*},\Psi_{20}^{*},\chi_{1\mathbf{k}}^{*},\chi_{2\mathbf{k}}^{*}\}$ yields a set of saddle-point equations Wang et al. (2024). Those derived from the derivative $\partial\Omega/\partial\chi_{\sigma\mathbf{k}}^{*}$ can be recast into the form of the BCS-like gap equation by introducing the Cooper-like pairing order parameter $\Delta_{\sigma}\equiv g_{\sigma}V^{-1}\sum_{\mathbf{k}\neq 0}\langle a_{\sigma\mathbf{k}}a_{\sigma,-\mathbf{k}}\rangle$. Using the pairing order parameter $\Delta_{\sigma}$ one can rewrite the ground state energy $\Omega$ in Eq. (8) as a function of only five unknowns: $\Omega=\Omega[\Psi_{10},\Psi_{20},\Delta_{1},\Delta_{2},\mu]$, whose first-order derivative with respect to the five parameters lead to four saddle-point equations plus one total particle number density constraint: $n=(|\Psi_{10}|^{2}+n_{1})+2(|\Psi_{20}|^{2}+n_{2})$ with $n_{\sigma}=V^{-1}\sum_{\mathbf{k}\neq 0}\langle a_{\sigma\mathbf{k}}^{\dagger}a_{\sigma\mathbf{k}}\rangle$.
We now numerically solve the five equations. Figures 3(a,b) plot the calculated chemical potentials near resonance at two different densities for both the wide and narrow resonances. This provides useful insight into the anomalous negative sign of the compressibility and the related condensate instability. From both figures we observe that the chemical potential $\mu$, which represents the average energy per atom, falls below the two energy levels corresponding to the bare atomic continuum threshold on the atomic side and the two-body molecular energy$/2$ on the molecular side. Furthermore, as the density $n$ decreases, in the near-resonance regime the chemical potential $\mu$ approaches the two energy levels from below. Consequently, the inverse compressibility $\kappa^{-1}=d\mu/dn$ must be negative near and on both sides of the resonance  [^7] In contrast to the near-resonance regime in Figs. 3(a,b), the chemical potential $\mu$ rises above one of the two energy levels when the detuning is far off the resonance. This is true for both the wide- and narrow-resonance cases (see Ref. sup for additional details). It implies that, in these far-off-resonance regimes, the chemical potential must approach these levels from above as the density vanishes $n\rightarrow 0$, consistent with the fact that the compressibility is positive in those detuning regimes. . The numerical results which were shown previously in Figs. 2(a,b) support this analysis; there we directly evaluated the derivative $d\mu/dn$ numerically, from which we constructed the ground-state stability phase diagrams.
Figure 3: Chemical potential, inverse compressibility and molecular scattering lengths. (a,b) Behavior of the chemical potential $\mu$ (blue solid and red dashed lines) versus detuning $\bar{\nu}$ at two different densities, $n$ and $n/2$ with $na_{\mathrm{bg}}^{3}=1.68\times 10^{-5}$ the same density value used in Figs. 2(c,d). “Atom continuum” threshold corresponds to $\mu=0$, and “$-E_{b}$” in (a) is the two-body but dressed molecular energy, which is replaced by its bare value $\bar{\nu}$ in (b). (c,d) Comparison of the corresponding inverse compressibility (blue solid lines) for the same density $na_{\mathrm{bg}}^{3}=1.68\times 10^{-5}$ with the atomic scattering length $a_{s}$ (red dashed lines) and the many-body molecular scattering length $a_{\mathrm{mm}}^{\mathrm{eff}}$ (brown dashed lines with squares), which is defined in Eq. (9). Both $a_{\mathrm{mm}}^{\mathrm{eff}}$ and $a_{s}$ are plotted in units of $a_{\mathrm{bg}}$.
Many-body effective scattering length — Armed with this understanding we now derive the many-body effective scattering length $a_{\mathrm{mm}}^{\mathrm{eff}}$ for the dressed molecules in terms of an effective interaction $g_{2}^{\mathrm{eff}}$:
$a_{\mathrm{mm}}^{\mathrm{eff}}(n)$ $=\frac{m_{2}}{4\pi\hbar^{2}}g_{2}^{\mathrm{eff}}(n)$ (9a)
with $\text{with }\quad g_{2}^{\mathrm{eff}}$ $=\frac{(\mathcal{Z}^{\mathrm{eff}})^{2}}{2}\frac{\partial^{4}\Omega[\Psi_{10},\Psi_{20},\Delta_{1},\Delta_{2},\mu]/V}{\partial\Psi_{20}^{2}\partial(\Psi_{20}^{*})^{2}}.$ (9b)
Here, $\mathcal{Z}^{\mathrm{eff}}$ is the fraction of the dressed-molecular condensate in the closed channel, defined by $\mathcal{Z}^{\mathrm{eff}}=|\Psi_{20}|^{2}/(|\Psi_{20}|^{2}+n_{1}/2)$  [^8] The factor $(\mathcal{Z}^{\mathrm{eff}})^{2}$ in Eq. (9b) is needed to reconcile two facts: (1) the many-body scattering length $a_{\mathrm{mm}}^{\mathrm{eff}}$, (as well as the two-body one $a_{\mathrm{mm}}$), is defined for a dressed molecule sup , which is an admixture of closed-channel molecules with open-channel atom pairs, and (2) the derivative in Eq. (9b) is taken with respect to the bare closed-channel molecular condensate amplitude $\Psi_{20}$. Although this factor is essentially equal to $1$ for a narrow resonance, it becomes important for wide resonances, particularly in deriving Eq. (2sup in the zero density limit. . The quartic derivative in Eq. (9b) can be carried out approximately sup  [^9] When performing derivatives in Eq. (9b) we set $\{\Psi_{10},\Delta_{1},\Delta_{2}\}$ at their saddle-point values. They are viewed as functions of $\{\Psi_{20},\mu\}$ only. In arriving at Eq. (10) we have also neglected terms that are higher order in the ratios $\{g_{1}n/E_{1\mathbf{k}=0},g_{2}n/E_{1\mathbf{k}=0},\alpha\sqrt{n}/E_{1\mathbf{k}=0}\}$. , leading to
$\frac{g_{2}^{\mathrm{eff}}}{(\mathcal{Z}^{\mathrm{eff}})^{2}}\approx\widetilde{\bar{g}}_{2}-\widetilde{\bar{\alpha}}^{4}\big{[}\frac{1}{V}\sum_{\mathbf{k}}\frac{1}{2E_{1\mathbf{k}}^{3}}-2g_{1}\big{(}\frac{1}{V}\sum_{\mathbf{k}}\frac{\widetilde{\epsilon}_{1\mathbf{k}}}{2E_{1\mathbf{k}}^{3}}\big{)}^{2}\big{]},$ (10)
where $E_{1\mathbf{k}}=\sqrt{\widetilde{\epsilon}_{1\mathbf{k}}^{2}-|\widetilde{\Delta}_{1}|^{2}}$ is the atomic Bogoliubov quasiparticle energy, with $\widetilde{\epsilon}_{1\mathbf{k}}=h_{1\mathbf{k}}+2g_{1}(|\Psi_{10}|^{2}+n_{1})$ and $\widetilde{\Delta}_{1}=\Delta_{1}+g_{1}\Psi_{10}^{2}-2\alpha\Psi_{20}$. In Eq. (10), $\widetilde{\bar{\alpha}}$ and $\widetilde{\bar{g}}_{2}$ are two interaction parameters, related to the Feshbach coupling $\bar{\alpha}$ and molecule-molecule interaction $\bar{g}_{2}$ by $\widetilde{\bar{\alpha}}=\sqrt{2}\alpha/(1+g_{1}V^{-1}\sum_{\mathbf{k}}1/2E_{1\mathbf{k}})$ and $\widetilde{\bar{g}}_{2}=g_{2}/(1+g_{2}V^{-1}\sum_{\mathbf{k}}1/2E_{2\mathbf{k}})$. It is important to note that Eq. (10) is applicable as long as the detuning $\bar{\nu}$ is smaller than and not too close to its QCP value $\bar{\nu}_{c}$ as we have already set $\Psi_{10}=0$. Evaluating Eqs. (9) and (10) leads to the $a_{\mathrm{mm}}^{\mathrm{eff}}$ plots in Figs. 2(c,d).
The overall minus sign associated with the term $\sum_{\mathbf{k}}1/2E_{1\mathbf{k}}^{3}$ in Eq. (10) should be noted. This term arises from contributions related to the scattering process shown in Fig. 1(b). It should be clear that the presence of $E_{1\mathbf{k}}$ in the denominator, reflects the fact that the excitations involved in the intermediate scattering state are atomic Bogoliubov quasiparticles.
In general, the $\mathbf{k}$ integral in Eq. (10) can not be done analytically. However, for the narrow resonance case and detuning near the QCP, Eqs. (10) and (9) can be further simplified to yield the simple analytical expression for the many-body scattering length $a_{\mathrm{mm}}^{\mathrm{eff}}$ sup presented in Eq. (4). In Figs. 3(c,d), we numerically evaluate $a_{\mathrm{mm}}^{\mathrm{eff}}$ for a generic detuning in both the narrow and wide resonance cases, and compare it with the numerically calculated inverse compressibility $\kappa^{-1}$   [^10] At large negative detuning $\bar{\nu}$, the many-body scattering length $a_{\mathrm{mm}}^{\mathrm{eff}}$ approaches its two-body background counterpart $a_{\mathrm{mm}}^{\mathrm{bg}}$.. Notably, the two, $a_{\mathrm{mm}}^{\mathrm{eff}}$ and $\kappa^{-1}$, show rather precise agreement, for both the narrow and wide resonances at detunings away from the immediate vicinity of the QCP  [^11] We note that the more extended unstable regime in Fig. 2(a) for a wide resonance is associated with the presence of a sizable number of atom Cooper pairs sup , which is also a consequence of the larger Feshbach coupling strength $\bar{\alpha}$. .
All of this allows us to understand why the compressibility in Fig. 2(a,b) behaves so differently when comparing the behavior of wide and narrow resonances. For a very narrow resonance, because of the small Feshbach coupling $\bar{\alpha}$, the factor $\widetilde{\bar{\alpha}}^{4}$ in Eq. (10), which is proportional to $\bar{\alpha}^{4}$, is very small. Consequently the effective scattering length $a_{\mathrm{mm}}^{\mathrm{eff}}$ becomes negative only when the atomic Bogoliubov quasiparticle energy gap, (contained in $E_{1\mathbf{k}=0}$), is sufficiently small. This insures that the detuning at which $a_{\mathrm{mm}}^{\mathrm{eff}}$ changes sign is sufficiently close to the QCP. By contrast, for a wide resonance this occurs when the gap $E_{1\mathbf{k}=0}$ is still large, which corresponds to a detuning well away from the QCP.
Zero density limit —The expressions for the two-body scattering length $a_{\mathrm{mm}}$, presented in Eqs. (2) and (3), were obtained from the zero-density limit of $a_{\mathrm{mm}}^{\mathrm{eff}}$ in Eq. (9), which leads to
$a_{\mathrm{mm}}=a_{\mathrm{mm}}^{\mathrm{bg}}-\frac{a_{\mathrm{mm}}^{\mathrm{bg}}}{k_{b}^{4}r_{*}\bar{a}^{3}}A_{1}-\frac{1}{k_{b}^{6}r_{*}^{2}\bar{a}^{3}}A_{2}+\frac{a_{\mathrm{bg}}}{k_{b}^{8}r_{*}^{2}\bar{a}^{6}}A_{3},$ (11)
where $r_{*}$ was defined in Eq. (6) and $k_{b}=\sqrt{mE_{b}}/\hbar$ is the detuning-dependent momentum corresponding to the molecular binding energy $E_{b}$. The three dimensionless and positive coefficients $\{A_{1},A_{2},A_{3}\}$ contain sub-leading dependences on $1/k_{b}$ sup , and their expressions are given in the Supplemental Materials sup . To arrive at Eqs. (2) and (3), we retain the background contribution $a_{\mathrm{mm}}^{\mathrm{bg}}$ in Eq. (11) along with the $A_{1}$ and $A_{2}$ terms, which are dominant for the wide and narrow resonances, respectively, near the boundary detuning $\bar{\nu}_{c,-}$ in Fig. 2.
Conclusions — In this Letter we have presented an in-depth study of the stability of molecular condensates in the context of Feshbach resonances of bosonic atoms. This has been a longstanding challenge for the field. By numerically calculating the compressibility we have not only constructed ground-state phase diagrams, but also derived concrete expressions for the two-body molecular scattering length $a_{\mathrm{mm}}$ and its many-body analogue $a_{\mathrm{mm}}^{\mathrm{eff}}$, given in Eqs. (2), (3), (4), and (11). These formulae, which through sign changes, allow one to infer the instability regions of molecular condensates, can and should be directly tested in future experiments. The concrete expressions for the molecular scattering lengths we derived make it possible to address a number of issues which were previously inaccessible; these include, among others, the equilibrium equation of state in the dilute molecular gas regime, the collective modes, the molecule-molecule scattering induced relaxation Petrov et al. (2005); Levinsen and Gurarie (2006); Levinsen and Petrov (2011), and the formation of the tetramer bound state Chin et al. (2005); Chen et al. (2023).
Acknowledgment — We thank P. Julienne and B. Zhao for helpful discussions. Z. W. and Q. C. are supported by the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0301904). This work is also supported by the National Science Foundation under Grant No. PHY-2409612 and PHY-2103542, and by the Air Force Office of Scientific Research under award number FA9550-21-1-0447. Z. Z. acknowledges the Bloch Postdoctoral Fellowship. We also acknowledge the University of Chicago’s Research Computing Center for their support of this work.